### v2

# Discrete Duality Symmetry in a $3D$ Field-Theoretic Model

Abstract: We demonstrate the discrete duality symmetry between the Abelian 1-form and 2-form basic gauge fields in the context of a three $(2+1)$-dimensional ($3D$) combined system of the field-theoretic model for the combined system of the free Abelian 1-from and 2-form gauge theories within the framework of Becchi-Rouet-Stora-Tyutin (BRST) formalism where the classical gauge-fixed Lagrangian density of this theory is generalized to its quantum counterpart as the BRST and co-BRST invariant Lagrangian density. We show clearly the existence of the off-shell nilpotent (co-)BRST symmetry transformations and establish their intimate connection through a set of underlying discrete duality symmetry transformations in our BRST-quantized theory. We provide the mathematical basis for the existence of the discrete duality symmetry transformations in our theory through the Hodge duality operator (that is defined on the $3D$ flat Minkowskian spacetime manifold). We briefly mention a bosonic symmetry transformation which is constructed from the anticommutator of the above off-shell nilpotent (co-)BRST symmetry transformations. We lay emphasis on the algebraic structures of the existing continuous and discrete duality symmetry transformations in our theory (where they are treated as operators).
PACS numbers: 11.15.-q; 11.30.-j; 11.10.Ef
Keywords: A $3D$ field-theoretic model; combined system of the Abelian 1-form and 2-form gauge theories; first-class constraints; (dual-)gauge symmetries; (anti-)BRST symmetries; (anti-)co-BRST symmetries; discrete duality symmetry transformations
1 Introduction
The study of the higher $p$-form ($p=2,3,4,...$) gauge theories***The most successful theory of THEP is the standard model of particle physics (SMPP) which is based on the interacting non-Abelian 1-form (i.e. $p=1$) gauge theory where there is a stunning degree of agreement between the theory and experiment. However, the conclusive experimental evidence of the masses of the neutrinos has compelled the modern-day practitioners of THEP to go beyond the domain of the validity of SMPP. The ideas behind the (super)string theories are very promising in this direction. has become important and essential due to the appearance of their basic fields in the quantum excitations of (super)strings which are the forefront areas of research activities in the realm of theoretical high energy physics (THEP). The modern developments in the domain of (super)string theories (see, e.g. [1-5] and references therein) have brought together, in an unprecedented manner, the active researchers in THEP and pure mathematics on a single intellectual platform where both kinds of researchers have benefited from the knowledge and creativity of each-other. In other words, there has been a notable convergence of ideas from the research activities in the realms of THEP and pure mathematics. Such kinds of confluence of ideas have been observed in our studies of the massless and Stückelberg-modified massive Abelian $p$-form (i.e. $p=1,2,3$) gauge theories (see, e.g. [6-9] and references therein) within the framework of Becchi-Rouet-Stora-Tyutin (BRST) formalism [10-13] where it has been established that, in the $D=2p$ (i.e. $D=2,4,6$) dimensions of spacetime, these theories turn out to be the field-theoretic examples for Hodge theory. The discrete and continuous symmetry transformations (and corresponding conserved charges) of these theories have been able to provide the physical realizations of the de Rham cohomological operators of differential geometry (see, e.g. [14-18]) at the algebraic level. In other words, in our research activities, it has been established that there is a convergence of ideas from the physical aspects of the BRST formalism and the mathematical ingredients related to the cohomological operators of differential geometry. It is crystal clear, from our present discussions, that our studies [6-9] on the above BRST-quantized Abelian theories have been confined only to the even dimensions (i.e. $D=2,4,6$) of the flat Minkowskian spacetime.
The central purpose of our present investigation is to focus on an odd dimensional field-theoretic example where we consider the combined system of a three $(2+1)$-dimensional ($3D$) free Abelian 1-form and 2-form gauge theories (within the framework of BRST formalism) to demonstrate that there is a set of discrete duality symmetries in this theory which connects the basic Abelian 1-form and 2-form gauge fields in a meaningful manner (i) at the level of the classically gauge-fixed Lagrangian density [cf. Eq. (12) below], and (ii) at the quantum level which is described by the (co-)BRST invariant Lagrangian density [cf. Eq. (21) below]. This observation plays a crucial role in providing the physical realizations of the algebraic structures of the cohomological operators of differential geometry. To be precise, these duality symmetry transformations connect the off-shell nilpotent BRST and co-BRST symmetry transformations that are respected by the BRST-invariant Lagrangian density of our theory [cf. Eq. (21) below]. The latter also respects a set of bosonic symmetry transformations that is obtained from the anticommutator of the off-shell nilpotent BRST and co-BRST symmetry transformations. One of the highlights of our present endeavor is the observation that the operator forms of the nilpotent (co-)BRST and non-nilpotent bosonic symmetry transformations obey an algebra that is reminiscent of the algebra that is satisfied by the de Rham cohomological operators of differential geometry.
The existence of the discrete duality symmetry transformations in our $3D$ theory is very important as they provide (i) a beautiful algebraic relationship between the off-shell nilpotent (co-)BRST symmetry transformations (cf. Sec. 4), and (ii) a direct connection between the fermionic (i.e. nilpotent) (co-)BRST symmetry transformations of our $3D$ theory (cf. Appendix A). In our recent couple of papers [19,20], we have not been able to obtain such kinds of relationships due to the fact that we were unaware of the existence of such kinds of beautiful discrete duality symmetry transformations. As a consequence, we could not provide the physical realization of the Hodge duality operator (in our earlier works [19,20]) which plays a key role in the algebraic structures that are satisfied by the de Rham cohomological operators of differential geometry. In our present endeavor, we have given a great deal of importance to the discrete duality symmetry transformations because they provide (i) a deep connection between the Abelian 1-form and 2-form basic gauge fields [cf. Eq. (3)], and (ii) the physical realization of the Hodge duality operator [14-18] the relationship: $\delta=\pm\,*\,d\,*$ between the (co-)exterior derivatives [i.e. $(\delta)d$].
One of the key upshots of our present endeavor is the observation that there is existence of the massless (i.e. $\Box\widetilde{\phi}=0,\;\Box\phi=0$) (pseudo-)scalar fields [i.e. $(\widetilde{\phi})\phi$] in our theory which carry the kinetic terms that are endowed with the opposite signs. However, despite this obvious distinct difference, they obey the massless Klein-Gordon equations of motion: $\Box\widetilde{\phi}=0,\;\Box\phi=0$. It is interesting to point that the fields with negative kinetic terms have become quite popular in the realm of the cosmological models of the Universe. In our theory, on the symmetry grounds alone, we observe that the pseudo-scalar field ($\widetilde{\phi}$) carries the negative kinetic term. Such kinds of fields are a set of possible candidates for (i) the “phantom” and/or “ghost” fields in the realm of the cyclic, bouncing and self-accelerated cosmological models of the Universe [21-25], and (ii) the dark energy (see, e.g. [26,27] and references therein) because these fields automatically lead to the creation of the negative pressure in the theory (which happens to be one of the characteristic features of dark energy). On the contrary, the pure scalar field ($\phi$), with the positive kinetic term, is a normal field which does not lead to the existence of the negative pressure in the theory. Hence, the existence of the pseudo-scalar field (in our theory) is very interesting in the current scenario of research activities that are related to(i) the realms of the cosmological models of the Universe, and (ii) the search for the possible dark energy candidate(s).
The theoretical materials of our present investigation are organized as follows. In Sec. 2, we provide a brief synopsis of the first-class constraints and ensuing classical gauge symmetry transformations for the combined system of the $3D$ free Abelian 1-form and 2-form gauge theories. In this section, we also discuss the gauge-fixed classical Lagrangian density and the continuous as well as discrete duality symmetry transformations associated with it. Our Sec. 3 deals with the elevation of the gauge-fixed classical Lagrangian density to its quantum counterpart as the (co-)BRST invariant Lagrangian density (where we show explicitly the existence of the nilpotent (co-)BRST symmetry transformations). The subject matter of our Sec. 4 is connected with the derivations of (i) a bosonic symmetry transformation which is constructed from the anticommutator of the (co-)BRST transformations, and (ii) the algebraic structures amongst the symmetry operators. The latter establish an intimate connection between the BRST and co-BRST transformations through the application of the discrete duality symmetry transformations. Finally, in our Sec. 5, we summarize our key results and point out the future perspective of our present investigation.
In our Appendix A, we establish a direct connection between the BRST and co-BRST symmetry transformations by exploiting the theoretical beauty and strength of the discrete duality symmetry transformations of our $3D$ theory.
2 Preliminaries: Constraint Analysis, (Dual-)Gauge and Discrete Duality Symmetry Transformations
We begin with the $3D$ field-theoretic model for the combined system of the free Abelian 1-form (i.e. $A^{(1)}=A_{\mu}\,dx^{\mu}$) and 2-form [i.e. $B^{(2)}=\frac{1}{2!}\,B_{\mu\nu}(dx^{\mu}\wedge dx^{\nu})$] gauge theories which is described by the following standard Lagrangian densityOur three $(2+1)$-dimensional ($3D$) Minkowskian spacetime is characterized by the flat metric tensor $\eta_{\mu\nu}=\text{diag}\,(+1,-1,-1)$ so that the dot product between two non-null Lorentz vectors $U_{\mu}$ and $V_{\mu}$ is defined as: $U\cdot V=\eta_{\mu\nu}\,U^{\mu}\,V^{\nu}\equiv U_{0}\,V_{0}-U_{i}\,V_{i}$ where the Greek indices $\mu,\nu,\sigma,...=0,1,2$ stand for the time and space directions and the Latin indices $i,j,k,...=1,2$ denote the space directions only. The $3D$ Levi-Civita tensor $\varepsilon_{\mu\nu\sigma}$ is chosen such that $\varepsilon_{012}=\varepsilon^{012}=+1$ and the standard relationships: $\varepsilon^{\mu\nu\sigma}\,\varepsilon_{\mu\nu\sigma}=+3!$, $\varepsilon^{\mu\nu\sigma}\,\varepsilon_{\mu\nu\eta}=+2!\,\delta^{\sigma}_{\eta}$, $\varepsilon^{\mu\nu\sigma}\,\varepsilon_{\mu\eta\kappa}=+1!\,(\delta^{\nu}_{\eta}\,\delta^{\sigma}_{\kappa}-\delta^{\nu}_{\kappa}\,\delta^{\sigma}_{\eta})$ are satisfied. We also adopt the convention of the derivative w.r.t. the antisymmetric tensor field as: $\big{(}\partial B_{\mu\nu}/\partial B_{\sigma\kappa}\big{)}=\frac{1}{2!}\,(\delta^{\sigma}_{\mu}\,\delta^{\kappa}_{\nu}-\delta^{\sigma}_{\nu}\,\delta^{\kappa}_{\mu}),etc.$
${\cal L}_{(0)}$ $=$ $-\dfrac{1}{4}\,F^{\mu\nu}\,F_{\mu\nu}+\dfrac{1}{12}\,H^{\mu\nu\sigma}\,H_{\mu\nu\sigma}\equiv-\dfrac{1}{2}\,\Big{(}\varepsilon^{\mu\nu\sigma}\,\partial_{\nu}A_{\sigma}\Big{)}^{2}+\dfrac{1}{2}\,\Big{(}\dfrac{1}{2}\varepsilon^{\mu\nu\sigma}\,\partial_{\mu}B_{\nu\sigma}\Big{)}^{2},$ (1)
where the field-strength tensors: $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $H_{\mu\nu\sigma}=\partial_{\mu}B_{\nu\sigma}+\partial_{\nu}B_{\sigma\mu}+\partial_{\sigma}B_{\mu\nu}$ are derived from the 2-form $F^{(2)}=d\,A^{(1)}=\frac{1}{2!}\,F_{\mu\nu}\big{(}dx^{\mu}\wedge dx^{\nu}\big{)}$ and 3-form $H^{(3)}=d\,B^{(2)}=\frac{1}{3!}\,H_{\mu\nu\sigma}\,\big{(}dx^{\mu}\wedge dx^{\nu}\wedge dx^{\sigma}\big{)}$, respectively. Here the symbol $d=\partial_{\mu}\,dx^{\mu}$ [with $d^{2}=\frac{1}{2!}\,\big{(}\partial_{\mu}\,\partial_{\nu}-\partial_{\nu}\,\partial_{\mu}\big{)}\big{(}dx^{\mu}\wedge dx^{\nu}\big{)}=0$] denotes the exterior derivative of differential geometry (see, e.g. [14-18]) and the above 1-form (i.e. $A^{(1)}$) and 2-form (i.e. $B^{(2)}$) define the Lorentz vector gauge field $A_{\mu}$ and the second-rank antisymmetric (i.e. $B_{\mu\nu}=B_{\nu\mu}$) tensor gauge field $B_{\mu\nu}$, respectively. It is the specific feature of the three $(2+1)$-dimensional $(3D)$ Minkowskian spacetime that the kinetic terms for the Abelian 1-form gauge field $A_{\mu}$ and the antisymmetric tensor gauge field $B_{\mu\nu}$ can be expressed in terms of the $3D$ Levi-Civita tensor and the derivative on them as we have expressed these terms in equation (1). One of the key reasons behind this observation is the fact that, in the $3D$ Minkowskian spacetime, we have: $\frac{1}{12}\,H^{\mu\nu\sigma}\,H_{\mu\nu\sigma}=\frac{1}{2}\,H^{012}\,H_{012}\equiv\frac{1}{2}\,(H_{012})^{2}$. The component $H_{012}$, however, can be written in its covariant form as: $H_{012}=\frac{1}{2}\,\varepsilon^{\mu\nu\sigma}\,\partial_{\mu}B_{\nu\sigma}$ which has been taken into account in our equation (1) to express the kinetic term for the Abelian 2-form gauge field. Similarly, it is straightforward to check that the $3D$ kinetic term for the Abelian 1-form field is: $-\frac{1}{4}\,F^{\mu\nu}\,F_{\mu\nu}=-\frac{1}{2}\,\big{(}\varepsilon^{\mu\nu\sigma}\partial_{\nu}A_{\sigma}\big{)}^{2}$. For our combined system of the $3D$ free Abelian 1-form and 2-form gauge theories, the infinitesimal, local and continuous gauge transformations $(\delta_{g})$ are well-known. These symmetry transformations for our $3D$ system are:
$\delta_{g}B_{\mu\nu}=-(\partial_{\mu}\Lambda_{\nu}-\partial_{\mu}\Lambda_{\nu}),\qquad\qquad\delta_{g}A_{\mu}=\partial_{\mu}\Lambda,$ (2)
where the Lorentz vector $\Lambda_{\mu}(x)$ and the Lorentz scalar $\Lambda(x)$ are the infinitesimal and local gauge symmetry transformation parameters. The above infinitesimal gauge symmetry transformations (2) are generated by the first-class constraints that exist on our theory.
To determine these first-class constraints, the first-step is to derive the canonical conjugate momenta (i.e. $\Pi^{\mu\mu}_{(B)},\;\Pi^{\mu}_{(A)}$ ) w.r.t. the basic gauge fields $B_{\mu\nu}$ and $A_{\mu}$, respectively. For the $3D$ version of the Lagrangian density (1), these momenta are:
$\Pi^{\mu\nu}_{(B)}$ $=$ $\dfrac{\partial{\cal L}_{(0)}}{\partial(\partial_{0}B_{\mu\nu})}\equiv\dfrac{1}{2}\,\varepsilon^{0\mu\nu}\,\Big{(}\dfrac{1}{2}\,\varepsilon_{\sigma\eta\kappa}\,\partial^{\sigma}B^{\eta\kappa}\Big{)}=\dfrac{1}{2}\,\varepsilon^{0\mu\nu}\,H_{012}\;\Longrightarrow\;\Pi^{0i}_{(B)}=\dfrac{1}{2}\,\varepsilon^{00i}\,H_{012}\approx 0,$
$\Pi^{\mu}_{(A)}$ $=$ $\dfrac{\partial{\cal L}_{(0)}}{\partial(\partial_{0}A_{\mu})}=-\varepsilon^{0\mu\sigma}\,\Big{(}\varepsilon_{\sigma\nu\kappa}\,\partial^{\nu}A^{\kappa}\Big{)}\equiv-F^{0\mu}\quad\Longrightarrow\quad\Pi^{0}_{(A)}=-F^{00}\approx 0.$ (3)
The above equation demonstrates that we have two sets of primary constraints (i.e. $\Pi^{0i}_{(B)}\approx 0,\;\Pi^{0}_{(B)}\approx 0$) on our theory where the symbol $\approx 0$ stands for weakly zero in Dirac’s notation (see, e.g. [28,29] for details). The requirements of the time-evolution invariance of these primary constraints lead to the determination of the secondary constraints. The following Euler-Lagrange (EL) equations of motion (EoMs)
$\partial_{\mu}\Big{(}\dfrac{1}{2}\,\varepsilon^{\mu\nu\sigma}\,H_{012}\Big{)}=0,\quad\qquad\partial_{\mu}F^{\mu\nu}=0,$ (4)
would lead to the time-evolution invariance of the primary constraints for the choices: $\nu=0,\;\sigma=i$ and $\nu=0$ in the above equations, respectively. These can be written as
$\partial_{0}\Big{(}\dfrac{1}{2}\,\varepsilon^{00i}\,H_{012}\Big{)}+\partial_{j}\Big{(}\dfrac{1}{2}\,\varepsilon^{j0i}\,H_{012}\Big{)}$ $=$ $0\;\Longrightarrow\;\partial_{0}\Pi^{0i}_{(B)}=\partial_{j}\Pi^{ji}_{(B)}\approx 0,$
$\partial_{0}F^{00}+\partial_{i}F^{i0}$ $=$ $0\;\Longrightarrow\;\partial_{0}\Pi^{0}_{(A)}=\partial_{i}\Pi^{i}_{(A)}\approx 0,\qquad$ (5)
where $\Pi^{ij}_{(B)}=\frac{1}{2}\,\varepsilon^{0ij}\,H_{012}$ and $\Pi^{i}_{(A)}=-F^{0i}$ are the space components of the conjugate momenta that have been defined in equation (2). It is obvious that we have four sets of constraints on our theory out of which two sets (i.e. $\Pi^{0i}_{(B)}\approx 0,\;\Pi^{0}_{(A)}\approx 0$) are the primary constraints and the rest of the two (i.e. $\partial_{j}\Pi^{ji}_{(B)}\approx 0,\;\partial_{i}\Pi^{i}_{(A)}\approx 0$) are the secondary constraints [28,29]. Since all these constraints are expressed in terms of the components of the conjugate momenta (2), all of them commute among themselves. Hence, they belong to the first-class variety according to Dirac’s prescription for the classification scheme of constraints [16,17]. These first-class constraints generate the infinitesimal gauge symmetry transformations (2) which we explain in the following paragraph.
According to Noether’s theorem, the existence of the continuous symmetry transformations always leads to the derivations of the Noether conserved currents and corresponding charges. Due to the infinitesimal gauge symmetry transformations in equation (2), we find the following expression for the Noether gauge current $J^{\mu}_{(G)}$ for our $3D$ field-theoretic combined system of the free Abelian 1-form and 2-form theory
$J^{\mu}_{(G)}=-F^{\mu\nu}\,\big{(}\partial_{\nu}\Lambda\big{)}-\dfrac{1}{2}\varepsilon^{\mu\nu\sigma}H_{012}\,\big{(}\partial_{\nu}\Lambda_{\sigma}-\partial_{\sigma}\Lambda_{\nu}\,\big{)},$ (6)
which is conserved (i.e. $\partial_{\mu}J^{\mu}_{(G)}=0$) due the EL-EoMs in (4) and the antisymmetric properties of the tensors: $F^{\mu\nu}$ and $\varepsilon^{\mu\nu\sigma}$. This conserved Noether gauge current [cf. Eq. (6)] leads to the definition of the conserved Noether charge $Q_{G}=\int d^{2}x\,J^{0}_{(G)}$ as follows:
$Q_{G}=\int d^{2}x\Big{[}-F^{0\nu}\big{(}\partial_{\nu}\Lambda\big{)}-\dfrac{1}{2}\varepsilon^{0\nu\sigma}H_{012}\big{(}\partial_{\nu}\Lambda_{\sigma}-\partial_{\sigma}\Lambda_{\nu}\big{)}\Big{]}.$ (7)
Since we have constraints (i.e. $-F^{00}\approx 0,\;\frac{1}{2}\,\varepsilon^{00i}\,H_{012}\approx 0$) on our theory, we have to be careful in expanding the r.h.s. of the above charge $Q_{G}$ so that these constraints are not strongly set equal to zero. In other words, we have the following explicit expression:
$Q_{G}$ $=$ $\int d^{2}x\Big{[}-F^{00}\big{(}\partial_{0}\Lambda\big{)}-F^{0i}\big{(}\partial_{i}\Lambda\big{)}-\dfrac{1}{2}\varepsilon^{00i}H_{012}\big{(}\partial_{0}\Lambda_{i}-\partial_{i}\Lambda_{0}\big{)}$ (8)
-- $\dfrac{1}{2}\varepsilon^{0ij}H_{012}\big{(}\partial_{i}\Lambda_{j}-\partial_{j}\Lambda_{i}\big{)}\Big{]}.$
Taking into account the primary constraints and the space components of the momenta that are present in equation (2), we recast the above charge in the following form:
$Q_{G}=\int d^{2}x\Big{[}\Pi^{0}_{(A)}\,\big{(}\partial_{0}\,\Lambda\big{)}+\Pi^{i}_{(A)}\big{(}\partial_{i}\Lambda\big{)}-\Pi^{0i}_{(B)}\big{(}\partial_{0}\Lambda_{i}-\partial_{i}\Lambda_{0}\big{)}-\Pi^{ij}_{(B)}\big{(}\partial_{i}\Lambda_{j}-\partial_{j}\Lambda_{i}\big{)}\Big{]}.$ (9)
Using the following non-zero equal-time commutatorsAt the classical level, we should use the Poisson-brackets instead of the canonical commutators. However, since we shall be discussing the BRST-quantized version of our present classical $3D$ theory, we shall stick with the canonical commutators which will be useful for our discussions at the quantum level.
$\big{[}B_{0i}(\vec{x},t),\,\Pi^{0j}_{(B)}(\vec{y},t)\big{]}$ $=$ $i\,\delta^{j}_{i}\,\delta^{(2)}(\vec{x}-\vec{y}),$
$\big{[}B_{ij}(\vec{x},t),\,\Pi^{kl}_{(B)}(\vec{y},t)\big{]}$ $=$ $\dfrac{i}{2}\big{(}\delta^{k}_{i}\delta^{l}_{j}-\delta^{k}_{j}\delta^{l}_{i}\big{)}\delta^{(2)}(\vec{x}-\vec{y}),$
$\big{[}A_{0}(\vec{x},t),\,\Pi^{0}_{(A)}(\vec{y},t)\big{]}$ $=$ $i\,\delta^{(2)}(\vec{x}-\vec{y}),$
$\big{[}A_{i}(\vec{x},t),\,\Pi^{j}_{(A)}(\vec{y},t)\big{]}$ $=$ $i\,\delta^{j}_{i}\,\delta^{(2)}(\vec{x}-\vec{y}),$ (10)
(in the natural units where we have: $\hbar=c=1$), it is straightforward to check that we obtain the infinitesimal gauge symmetry transformations (2) from the conserved Noether charge $Q_{G}$ [cf. Eq. (9)] if we use the standard relationship between the continuous symmetry transformation $(\delta_{g})$ and its generator $Q_{G}$ as: $\delta_{g}\Phi(\vec{x},t)=-i\,\big{[}\Phi(\vec{x},t),\;Q_{G}\big{]}$ where the generic field $\Phi(\vec{x},t)\equiv A_{0}(\vec{x},t),\,A_{i}(\vec{x},t),\,B_{0i}(\vec{x},t),\,B_{ij}(\vec{x},t)$. The above expression for the conserved Noether charge (9) can be re-written in terms of the first-class constraints [cf. Eqs. (2),(2)] as the generator $G$ of the infinitesimal gauge symmetry transformations (2) by exploiting the theoretical strength of the Gauss divergence theorem as:
$Q_{G}\to G$ $=$ $\int d^{2}x\Big{[}\Pi^{0}_{(A)}\big{(}\partial_{0}\Lambda\big{)}-\big{(}\partial_{i}\Pi^{i}_{(A)}\big{)}\Lambda-\Pi^{0i}_{(B)}\big{(}\partial_{0}\Lambda_{i}-\partial_{i}\Lambda_{0}\big{)}$ (11)
$+$ $\big{(}\partial_{i}\Pi^{ij}_{(B)}\big{)}\Lambda_{j}+\big{(}\partial_{j}\Pi^{ji}_{(B)}\big{)}\Lambda_{i}\Big{]},$
which matches with the general expression for the generator $G$, in terms of the first-class constraints, that has been written in a nice paper (see, e.g. [30] for details).
For the quantization of our combined system of the free Abelian 1-form and 2-form gauge theories, we have to incorporate the gauge-fixing terms [i.e. ${\cal L}_{(gf)}$] into the starting Lagrangian density [cf. Eq. (1)]. The ensuing gauge-fixed Lagrangian density is:
${\cal L}_{(0)}+{\cal L}_{(gf)}=-\dfrac{1}{2}\Big{(}\varepsilon^{\mu\nu\sigma}\,\partial_{\nu}A_{\sigma}\Big{)}^{2}-\dfrac{1}{2}\big{(}\partial\cdot A\big{)}^{2}+\dfrac{1}{2}\Big{(}\dfrac{1}{2}\varepsilon^{\mu\nu\sigma}\,\partial_{\mu}B_{\nu\sigma}\Big{)}^{2}+\dfrac{1}{2}\big{(}\partial^{\nu}B_{\nu\mu}\big{)}^{2}.$ (12)
The gauge-fixing terms in the above owe their origin to the co-exterior derivative $\delta=\pm\,*\,d\,*$ of the differential geometry [14-18]. For instance, it can be readily checked that: $\delta\,A^{(1)}=\pm\,*\,d\,*\big{(}A_{\mu}\,dx^{\mu}\big{)}\equiv\pm\big{(}\partial\cdot A\big{)}$ and $\delta\,B^{(2)}=\pm\,*\,d\,*\big{[}\frac{1}{2!}\,B_{\mu\nu}\,\big{(}dx^{\mu}\wedge dx^{\nu}\big{)}\big{]}\equiv\mp\big{(}\partial^{\nu}B_{\nu\mu}\big{)}dx^{\mu}$ where * is the Hodge duality operator on our chosen $3D$ Minkowskian spacetime manifold. Under the infinitesimal local gauge symmetry transformations (2), we obtain the following:
$\delta_{g}\big{[}{\cal L}_{(0)}+{\cal L}_{(gf)}\big{]}=\partial_{\mu}\Big{[}\big{(}\partial_{\nu}B^{\nu\mu}\big{)}\big{(}\partial\cdot\Lambda\bigl{)}\Big{]}-\big{(}\partial\cdot A\big{)}\,\Box\Lambda-\big{(}\partial_{\nu}B^{\nu\mu}\big{)}\,\Box\Lambda_{\mu}.$ (13)
The above equation demonstrates that if we impose the conditions: $\Box\Lambda=0$ and $\Box\Lambda_{\mu}=0$ on the Lorentz scalar and vector gauge transformation parameters, we have the gauge invariance in the theory because the gauge-fixed Lagrangian density (12) transforms to a total spacetime derivative (thereby rendering the corresponding action integral invariant). We have another infinitesimal, local and continuous symmetry in the theory which we call as the dual-gauge symmetry transformations $(\delta_{dg})$, under which, we have the following transformations for the Abelian 1-form and 2-form basic gauge fields, namely;
$\delta_{dg}A_{\mu}=-\,\varepsilon_{\mu\nu\sigma}\partial^{\nu}\Sigma^{\sigma},\qquad\qquad\delta_{dg}B_{\mu\nu}=+\,\varepsilon_{\mu\nu\sigma}\partial^{\sigma}\Sigma,$ (14)
where the Lorentz axial-vector $\Sigma_{\mu}(x)$ and pseudo-scalar $\Sigma(x)$ are the infinitesimal local dual-gauge symmetry transformation parameters (so that the parity symmetry could be maintained at every stage of the above transformations). Under the above infinitesimal dual-gauge symmetry transformations, the gauge-fixed Lagrangian density transforms to:
$\delta_{dg}\big{[}{\cal L}_{(0)}+{\cal L}_{(gf)}\big{]}=\partial_{\mu}\Big{[}\varepsilon^{\mu\nu\sigma}\,\partial_{\nu}A_{\sigma}\big{(}\partial\cdot\Sigma\big{)}\Big{]}+\frac{1}{2}\,\Big{(}\varepsilon^{\mu\nu\sigma}\,\partial_{\mu}B_{\nu\sigma}\Big{)}\Box\Sigma-\Big{(}\varepsilon^{\mu\nu\sigma}\,\partial_{\nu}A_{\sigma}\Big{)}\Box\Sigma_{\mu}.$ (15)
The above equation demonstrates that if we choose the infinitesimal dual-gauge transformation parameters such that the restrictions: $\Box\Sigma_{\mu}=0$ and $\Box\Sigma=0$ are satisfied together, we have a perfect dual-gauge symmetry in the theory because the gauge-fixed Lagrangian density transforms to a total spacetime derivative (under the above conditions). In addition to the infinitesimal (i) gauge symmetry transformations (2), and (ii) dual-gauge symmetry transformations (14) (with similar kinds of restrictions on the transformation parameters: $\Box\Lambda=0,\;\Box\Lambda_{\mu}=0,\;\Box\Sigma=0,\;\Box\Sigma_{\mu}=0$), we have the following perfect discrete duality symmetry transformations§§§Mathematical origin for the existence of the perfect duality symmetry transformations (16) for the $3D$ combined system of the free Abelian 1-form and 2-form gauge theories, seems to be hidden in the Hodge duality * operation on the Abelian 2-form and 1-form, respectively. For instance, we note that: $*\,B^{(2)}=\big{[}\frac{1}{2}\,\varepsilon_{\mu\nu\sigma}\,B^{\nu\sigma}\big{]}\,dx^{\mu}\sim A_{\mu}\,dx^{\mu}$ which demonstrate that, for our $3D$ theory, the Hodge dual of 2-form is the 1-form. In a similar manner, we observe that the Hodge dual of the 1-form: $*\,A^{(1)}=\frac{1}{2}\,\big{[}\varepsilon_{\mu\nu\sigma}\,A^{\sigma}\big{]}\,(dx^{\mu}\wedge dx^{\nu})\sim\frac{1}{2}\,B_{\mu\nu}\,(dx^{\mu}\wedge dx^{\nu})$ is the 2-form. These observations are the reasons behind the numeral factors that appear in the discrete duality transformations (16) modulo the overall factors of ±$\pm\,i$.:
$A_{\mu}\longrightarrow\,\mp\,\dfrac{i}{2}\,\varepsilon_{\mu\nu\sigma}\,B^{\nu\sigma},\quad\qquad B_{\mu\nu}\longrightarrow\,\pm\,i\,\varepsilon_{\mu\nu\sigma}\,A^{\sigma}.$ (16)
This is due to the fact that we note the following
$-\,\dfrac{1}{2}\,\big{(}\partial\cdot A\big{)}^{2}\longleftrightarrow+\,\dfrac{1}{2}\,\Big{(}\dfrac{1}{2}\varepsilon^{\mu\nu\sigma}\,\partial_{\mu}B_{\nu\sigma}\Big{)}^{2},\qquad+\,\dfrac{1}{2}\,\big{(}\partial^{\nu}B_{\nu\mu}\big{)}^{2}\longleftrightarrow-\,\dfrac{1}{2}\,\Big{(}\varepsilon^{\mu\nu\sigma}\,\partial_{\nu}A_{\sigma}\Big{)}^{2},$ (17)
where (i) the gauge-fixing term of the Abelian 1-form field interchanges with the kinetic term of the Abelian 2-form field, and (ii) the gauge-fixing term of the Abelian 2-form field interchanges with the kinetic term of the Abelian 1-form field. This observation is very crucial as it will play a key role at the quantum level when we shall discuss various kinds of continuous and discrete symmetries of our $3D$ system (as the BRST-quantized theory).
We conclude this section with the following remarks. First, under the gauge symmetry transformations (2), the kinetic terms (owing their origin to the exterior derivative of differential geometry) remain invariant. Second, the gauge-fixing terms (owing their existence to the co-exterior derivative of differential geometry) are found to be invariant under the dual-gauge symmetry transformations. Finally, similar kinds of outside restrictions: $\Box\Sigma=0,\;\Box\Sigma_{\mu}=0,\;\Box\Lambda=0,\;\Box\Lambda_{\mu}=0$ on the (dual-)gauge symmetry transformation parameters at the classical level will be taken care of by the (anti-)ghost fields at the quantum level within the framework of BRST formalism (without any kinds of outside restrictions).
3 BRST and co-BRST Invariant Lagrangian Density
The gauge-fixed Lagrangian density (12) for our $3D$ combined system of the free Abelian 1-form and 2-form gauge theories can be linearized, in its most general form in two different ways, as follows
(,${\cal L}^{(1)}_{({\cal B},B)}$ $=$ $\dfrac{1}{2}\,{\cal B}^{\mu}{\cal B}_{\mu}-{\cal B}^{\mu}\Big{(}\varepsilon_{\mu\nu\sigma}\partial^{\nu}A^{\sigma}-\dfrac{1}{2}\,\partial_{\mu}\widetilde{\phi}\Big{)}-B\,(\partial\cdot A)+\dfrac{B^{2}}{2}$ (18)
$+$ ${\cal B}\Big{(}\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma}\,\partial^{\mu}B^{\nu\sigma}\Big{)}-\dfrac{{\cal B}^{2}}{2}+B^{\mu}\Big{(}\partial^{\nu}B_{\nu\mu}-\dfrac{1}{2}\,\partial_{\mu}\phi\Big{)}-\dfrac{1}{2}\,B^{\mu}B_{\mu},$
[…]
$+$ ${\cal B}\Big{(}\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma}\,\partial^{\mu}B^{\nu\sigma}\Big{)}-\dfrac{{\cal B}^{2}}{2}-{\bar{B}}^{\mu}\Big{(}\partial^{\nu}B_{\nu\mu}+\dfrac{1}{2}\,\partial_{\mu}\phi\Big{)}-\dfrac{1}{2}\,{\bar{B}}^{\mu}{\bar{B}}_{\mu}.$
where the subscripts on the above Lagrangian densities denote the short forms of the Nakanishi-Lautrup auxiliary fields that have been invoked to linearize the kinetic and gauge-fixing terms of the combined system of the free Abelian 2-form and 1-form gauge theories. To be more specific, we note that, in the Lagrangian density (18), a pair of auxiliary fields (${\cal B}_{\mu},\,B$) have been invoked to linearize the kinetic and gauge-fixing terms of the $3D$ Abelian 1-form gauge theory, respectively. On the other hand, another pair of auxiliary fields (${\cal B},\,B_{\mu}$) have been incorporated into the theory to linearize the kinetic and gauge-fixing terms of the 3D Abelian 2-form theory, respectively. In exactly similar fashion, we can specify the the concise form of the subscript on the Lagrangian density (19) where the replacements for the pairs of auxiliary fields are: $({\cal B}_{\mu},\,B)\Rightarrow(\bar{\cal B}_{\mu},\,B)$ and $({\cal B},\,B_{\mu})\Rightarrow({\cal B},\,\bar{B}_{\mu})$, respectively. In the above Lagrangian densities (18) and (19), an additional pair of dynamical fields ($\widetilde{\phi},\;\phi$) have been added at appropriate places with proper mass dimensions in natural units (where $\hbar=c=1$). These fields are nothing but the pseudo-scalar and scalar fields, respectively. The above linearizations are general in the sense that we have taken into account both the signs that can be associated with the fields ($\widetilde{\phi},\;\phi$). In other words, a close look at (18) and (19) demonstrates that both the Lagrangian densities can be exchanged with each-other by taking into account the replacements: $(\widetilde{\phi},\;\phi)\leftrightarrow(-\,\widetilde{\phi},\;-\,\phi),\;({\cal B}_{\mu},\,B)\leftrightarrow(\bar{\cal B}_{\mu},\,B),\;({\cal B},\,B_{\mu})\leftrightarrow({\cal B},\,\bar{B}_{\mu})$. At this juncture, it is interesting to point out that the discrete duality symmetry transformations (16), existing for the gauge-fixed Lagrangian density (12), can be generalized into the following transformations:
$A_{\mu}\longrightarrow\,\mp\,\dfrac{i}{2}\,\varepsilon_{\mu\nu\sigma}\,B^{\nu\sigma},\quad B_{\mu\nu}\longrightarrow\,\pm i\,\varepsilon_{\mu\nu\sigma}\,A^{\sigma},\quad B\longrightarrow\,\mp\,i{\cal B},\qquad{\cal B}\longrightarrow\,\pm\,iB,$
$B_{\mu}\longrightarrow\,\mp\,i\,{\cal B}_{\mu},\qquad{\cal B}_{\mu}\longrightarrow\,\;\pm\,i\,B_{\mu},\qquad\bar{B}_{\mu}\longrightarrow\,\mp\,i\bar{\cal B}_{\mu},\qquad\bar{\cal B}_{\mu}\longrightarrow\,\pm\,i\bar{B}_{\mu},$
$\phi\longrightarrow\,\mp\,i\,\widetilde{\phi},\quad\widetilde{\phi}\longrightarrow\,\pm\,i\phi.$ (20)
The above discrete duality symmetry transformations are respected by the linearized versions of the Lagrangian densities (18) and (19).
For the sake of brevity, in our present endeavor, we consider only (i) the off-shell nilpotent (co-)BRST symmetry transformations which are the generalizations of the classical (dual-)gauge symmetry transformations (14) and (2) to the quantum level, and (ii) the (co-)BRST invariant Lagrangian density ${\cal L}_{B}$ which is the generalization of the classical linearized version of the Lagrangian density (,${\cal L}^{(1)}_{({\cal B},B)}$ [cf. Eq. (18)] to its quantum level. In other words, we promote the Lagrangian density (18) to incorporate into itself the Faddeev-Popov (FP) ghost terms corresponding to the $3D$ field-theoretic model of the combined system of the Abelian 1-form and 2-form gauge theories as follows
${\cal L}_{B}$ $=$ $\dfrac{1}{2}\,{\cal B}^{\mu}{\cal B}_{\mu}-{\cal B}^{\mu}\Big{(}\varepsilon_{\mu\nu\sigma}\partial^{\nu}A^{\sigma}-\dfrac{1}{2}\,\partial_{\mu}\widetilde{\phi}\Big{)}-B\,(\partial\cdot A)+\dfrac{B^{2}}{2}$ (21)
$+$ ${\cal B}\Big{(}\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma}\,\partial^{\mu}B^{\nu\sigma}\Big{)}-\dfrac{{\cal B}^{2}}{2}+B^{\mu}\Big{(}\partial^{\nu}B_{\nu\mu}-\dfrac{1}{2}\,\partial_{\mu}\phi\Big{)}-\dfrac{1}{2}\,B^{\mu}B_{\mu}$
-- $\big{(}\partial^{\mu}\bar{C}^{\nu}-\partial^{\nu}\bar{C}^{\mu}\big{)}\big{(}\partial_{\mu}C_{\nu}\big{)}-\dfrac{1}{2}\Big{(}\partial\cdot C-\dfrac{\lambda}{4}\Big{)}\rho-\dfrac{1}{2}\Big{(}\partial\cdot\bar{C}+\dfrac{\rho}{4}\Big{)}\lambda$
-- $\dfrac{1}{2}\partial^{\mu}\bar{\beta}\,\partial_{\mu}\beta-\partial^{\mu}\bar{C}\,\partial_{\mu}C,$
where the FP-terms, corresponding to the free massless Abelian 2-form theory, have been systematically derived in our earlier work [5]. In this work, we have considered the BRST-quantized version of the $4D$ Stückelberg-modified massive Abelian 2-form theory and the FP-ghost term (i.e. $-\,\partial^{\mu}\bar{C}\,\partial_{\mu}C$) for the free Abelian 1-form theory is the standard book material. In the above Lagrangian density (21), the Lorentz vector fermionic (i.e. $C_{\mu}^{2}=\bar{C}_{\mu}^{2}=0,\,C_{\mu}\,\bar{C}_{\nu}+\bar{C}_{\nu}\,C_{\mu}=0,\,C_{\mu}\,C_{\nu}+C_{\nu}\,C_{\mu}=0,$ etc.) (anti-)ghost fields $(\bar{C}_{\mu})C_{\mu}$ are the generalizations of the gauge transformations parameter $\Lambda_{\mu}$ [cf. Eq. (2)] and they carry the ghost numbers $(-1)+1$, respectively, and the bosonic (anti-)ghost fields $(\bar{\beta})\beta$ are endowed with the ghost numbers $(-2)+2$. Another pair of the fermionic (i.e. $C^{2}=\bar{C}^{2}=0,\,C\,\bar{C}+\bar{C}\,C=0$) (anti-)ghost fields $(\bar{C})C$ carry the ghost numbers $(-1)+1$ and they are the generalizations of the Lorentz scalar gauge transformation parameter $\Lambda$ [cf. Eq. (2)] and the auxiliary (anti-)ghost fields $(\rho)\lambda$ are endowed with ghost numbers $(-1)+1$, respectively. These (anti-)ghost fields are required to maintain the sacrosanct property of unitarity at any arbitrary order of perturbative computations. The other symbols in (21) carry the same meaning as explained in the context of the Lagrangian density in (18).
The following infinitesimal, continuous and off-shell nilpotent (i.e. $s_{d}^{2}=0,\;s_{b}^{2}=0$) (co-)BRST transformations $(s_{d})s_{b}$, namely;
$s_{d}B_{\mu\nu}=\varepsilon_{\mu\nu\sigma}\,\partial^{\sigma}\bar{C},\qquad\quad s_{d}A_{\mu}=-\varepsilon_{\mu\nu\sigma}\,\partial^{\nu}\bar{C}^{\sigma},\quad\qquad s_{d}\bar{C}_{\mu}=-\partial_{\mu}\bar{\beta},$
$s_{d}C_{\mu}=-{\cal B}_{\mu},\qquad\quad s_{d}C={\cal B},\qquad\quad s_{d}\beta=-\lambda,\qquad\quad s_{d}\widetilde{\phi}=\rho,$
$s_{d}\big{[}B_{\mu},\,{\cal B}_{\mu},\,B,\,{\cal B},\phi,\bar{\beta},\,\bar{C},\,\rho,\lambda\big{]}=0,\qquad$ (22)
$s_{b}B_{\mu\nu}=-\big{(}\partial_{\mu}C_{\nu}-\partial_{\nu}C_{\mu}\big{)},\qquad\quad s_{b}A_{\mu}=\partial_{\mu}C,\qquad\quad s_{b}C_{\mu}=-\partial_{\mu}\beta,$
$s_{b}\bar{C}_{\mu}=B_{\mu},\qquad\quad s_{b}\bar{C}=B,\qquad\quad s_{b}\bar{\beta}=-\rho,\qquad\quad s_{b}\phi=\lambda,$
$s_{b}\big{[}B_{\mu},\,{\cal B}_{\mu},\,B,\,{\cal B},\widetilde{\phi},\beta,\,C,\,\rho,\,\lambda\big{]}=0,\qquad$ (23)
are the symmetry transformationsThese off-shell nilpotent symmetry transformations are distinctly different from the corresponding transformations in our very recent publications [19,20]. of the action integral: $S=\int d^{3}x\,{\cal L}_{B}$ because the Lagrangian density ${\cal L}_{B}$ [cf. Eq. (21)] transforms to the total spacetime derivatives under the above nilpotent (co-)BRST [i.e. (dual-)BRST] symmetry transformations as follows:
$s_{d}{\cal L}_{B}=-\partial_{\mu}\Big{[}\big{(}\partial^{\mu}\bar{C}^{\nu}-\partial^{\nu}\bar{C}^{\mu}\big{)}\,{\cal B}_{\nu}-{\cal B}\,\partial^{\mu}\bar{C}-\frac{1}{2}\,\rho\,{\cal B}^{\mu}-\frac{1}{2}\,\lambda\,\partial^{\mu}\bar{\beta}\Big{]},$
$s_{b}\,{\cal L}_{B}=-\partial_{\mu}\Big{[}\big{(}\partial^{\mu}C^{\nu}-\partial^{\nu}C^{\mu}\big{)}\,B_{\nu}+B\,\partial^{\mu}C+\frac{1}{2}\,\lambda\,B^{\mu}-\dfrac{1}{2}\,\rho\,\partial^{\mu}\beta\Big{]}.$ (24)
In addition to the above off-shell nilpotent (co-)BRST symmetry transformations, the Lagrangian density (21) respects a set of discrete duality symmetry transformations in the ghost-sector which is listed as follows:
$C_{\mu}\,\longrightarrow\,\pm\,i\,\bar{C}_{\mu},\qquad\quad\bar{C}_{\mu}\,\longrightarrow\,\pm\,i\,C_{\mu},\qquad\quad C\,\longrightarrow\,\mp\,i\,\bar{C},$
$\bar{C}\,\longrightarrow\,\mp\,i\,C,\qquad\quad\quad\beta\,\longrightarrow\,\pm\,i\,\bar{\beta},\qquad\quad\quad\bar{\beta}\,\longrightarrow\,\mp\,i\,\beta,$
$\rho\,\longrightarrow\,\mp\,i\,\lambda,~{}\qquad\quad\quad\lambda\,\longrightarrow\,\mp\,i\,\rho.$ (25)
Thus, it is clear that the total (co-)BRST invariant Lagrangian density (21) respects the discrete duality symmetry transformations in the non-ghost sector [cf. Eq. (3)] as well as in the ghost-sector [cf. Eq. (3)]. To sum-up, the Lagrangian density (21) respects the infinitesimal, continuous and nilpotent symmetries [cf. Eqs. (3),(3)] and a set of discrete duality symmetry transformations. The latter is the sum of the discrete symmetry transformations that have been listed in (3) and (3)). We call these discrete symmetry transformations as the duality symmetry transformations because the basic 1-form and 2-form fields are connected with each-other [cf. Eq. (3)] and they are dual to each-other on a flat $3D$ Minkowskian spacetime manifold [cf. footnote below Eq. (16)].
4 Bosonic Symmetry and Algebraic Structures
It is straightforward to note, from the equations (3) and (3), that the anticommutator (i.e. $\{s_{b},\;s_{d}\}$) between them is not equal to zero. In fact, this anticommutator defines a bosonic symmetry (i.e. $s_{\omega}=\{s_{b},\;s_{d}\}$), under which, the fields of (21) transform as:
$s_{\omega}B_{\mu\nu}=\varepsilon_{\mu\nu\sigma}\,\partial^{\sigma}B+\big{(}\partial_{\mu}{\cal B}_{\nu}-\partial_{\nu}{\cal B}_{\mu}\big{)},\qquad s_{\omega}C_{\mu}=\partial_{\mu}\lambda,$
$s_{\omega}A_{\mu}=-\varepsilon_{\mu\nu\sigma}\,\partial^{\nu}B^{\sigma}+\partial_{\mu}{\cal B},\qquad s_{\omega}\bar{C}_{\mu}=\partial_{\mu}\rho,$
$s_{\omega}\big{[}B_{\mu},\,{\cal B}_{\mu},\,B,\,{\cal B},\,\phi,\,\widetilde{\phi},\,\beta,\,\bar{\beta},\rho,\,\lambda,\,C,\,\bar{C}\big{]}=0.$ (26)
At this stage, is worthwhile to mention that under the above bosonic symmetry transformations, the (anti-)ghost fields either do not transform at all or they transform up to the $U(1)$ gauge symmetry-type transformations. Furthermore, it is very interesting to point out that, in their operator forms, the nilpotent (co-)BRST symmetry transformations and the bosonic symmetry transformation obey the following algebra, namely;
$s_{b}^{2}=0,\quad\qquad s_{d}^{2}=0,\quad\qquad s_{\omega}=\big{\{}s_{b},\;s_{d}\big{\}}\equiv\big{(}s_{b}+s_{d}\big{)}^{2},$
$\big{[}s_{\omega},\;s_{b}\big{]}=0,\quad\qquad\big{[}s_{\omega},\;s_{d}\big{]}=0,\quad\qquad\big{\{}s_{b},\;s_{d}\big{\}}\neq 0,$ (27)
which establishes the fact that the bosonic symmetry transformation, in its operator form, commutes with the off-shell nilpotent (co-)BRST symmetry transformation operators. The algebra (4) is reminiscent of the following algebra obeyed by a set of three de Rham cohomological operators of differential geometry [14-18]
$d^{2}=0,\quad\qquad\delta^{2}=0,\quad\qquad\Delta=\big{\{}d,\;\delta\big{\}}\equiv\big{(}d+\delta\big{)}^{2},$
$\big{[}\Delta,\;d\big{]}=0,\quad\qquad\big{[}\Delta,\;\delta\big{]}=0,\quad\qquad\big{\{}d,\;\delta\big{\}}\neq 0,$ (28)
where $d$ (with $d^{2}=0$) is the exterior derivative, $\delta=\pm\,*\,d\,*$ (with $\delta^{2}=0$) is the co-exterior (or dual-exterior) derivative and $\Delta=(d+\delta)^{2}$ is the Laplacian operator (which is always positive definite). Here the symbol * stands for the Hodge duality operator on the compact spacetime manifold on which the cohomological operators are defined.
We end this short section with a couple of useful remarks. First of all, at the algebraic level, it is very tempting to appreciate the similarity between the above algebraic relationships in (4) and (4) which allow us to state that there is a one-to-one mappingWe have obtained the one-to-one mapping because we have considered only the Lagrangian density (21) which is the generalization of the linearized classical version of the gauge-fixed Lagrangian density (18) to its quantum level. If we had considered the quantum version of the linearized classical version of the gauge-fixed Lagrangian density (19) together with the above generalization, we would have ended up with the two-to-one mapping between the symmetry transformation operators and the de Rham cohomological operators of differential geometry as we have obtained in our earlier works [6-9]. between the symmetry transformation operators ($s_{b},\;s_{d},\;s_{\omega}$) and the de Rham cohomological operators ($d,\;\delta,\;\Delta$). To corroborate this claim, we have to recall our key observations in Secs. 2 and 3 where we have established that, under the (dual-)gauge and (co-)BRST symmetry transformations, the (i) gauge-fixing terms (owing their origin to the co-exterior derivative), and (ii) kinetic terms (owing their existence to the exterior derivative) for the gauge fields remain invariant, respectively. Hence, the above identification: $(s_{b},\;s_{d},\;s_{\omega})\Leftrightarrow(d,\;\delta,\;\Delta)$ is realistic at the algebraic level. Second, the (co-)exterior derivatives $(\delta)d$, as pointed out earlier, are connected by the relationship: $\delta=\pm\,*\,d\,*$. Within the framework of BRST formalism, this relationship is realized in terms of the nilpotent (co-)BRST transformation operators and the discrete transformation operators [cf. Eqs. (3),(3)] as
$s_{d}\,\Phi=\pm\,*\,s_{b}\,*\,\Phi,\qquad\Phi=B_{\mu\nu},A_{\mu},B_{\mu},{\cal B}_{\mu},\bar{C}_{\mu},C_{\mu},\phi,\widetilde{\phi},B,{\cal B},\bar{\beta},\beta,\bar{C},C,\rho,\lambda,$ (29)
where the symbol * stands for the discrete duality symmetry transformations. Thus, we observe that the interplay between the continuous and discrete symmetry operators of our theory provide the physical realization of the relationship: $\delta=\pm\,*\,d\,*$. In the above equation (29), as is obvious, the generic field of the Lagrangian density (21) has been denoted by the field $\Phi$. The ($\pm$) sign, on the r.h.s. of the above equation (29), is determined and dictated by a couple of successive operations of the discrete duality symmetry transformation operator on the generic field $\Phi$ of the Lagrangian density (21) as follows [31]:
$*\;\big{(}*\;\Phi\big{)}=\pm\,\Phi.$ (30)
It is an elementary exercise to note that, for all the bosonic fields of our (co-)BRST invariant Lagrangian density (21), we have the plus sign on the r.h.s. of (30) and, hence, in the equation (29), too. For instance, it is straightforward to check that: $*\,(*\,B_{\mu\nu})=+\,B_{\mu\nu}$. As a consequence, we observe that: $s_{d}B_{\mu\nu}=+\,*\,s_{b}\,*\,B_{\mu\nu}$ is true. On the other hand, the signs on the r.h.s. of (30) and (29) are negative when take into account the fermionic fields of the (co-)BRST invariant Lagrangian density (21). For instance, we can readily check that: $*\,(*\,C_{\mu})=-\,C_{\mu}$. As a consequence, we have the algebraic relationship between the off-shell nilpotent (co-)BRST symmetry transformations [cf. Eqs. (3),(3)] as: $s_{d}C_{\mu}=+\,*\,s_{b}\,*\,C_{\mu}$ where the symbol * stands for the set of discrete duality symmetry transformation operators that have been listed in (3) and (3).
5 Conclusions
In our present investigation, we have taken into account the combined system of the free $3D$ Abelian 1-form and 2-form gauge theories together and shown the existence of (i) the nilpotent BRST symmetry transformations, (ii) the nilpotent co-BRST symmetry transformations, and (iii) a bosonic symmetry transformation which is constructed from the anticommutator of the BRST and co-BRST symmetry transformations. Under the BRST symmetry transformations, the Lagrangian densities of the free Abelian 1-form and 2-form gauge theories remain invariant, separately and independently. However, for the validity of the co-BRST symmetry invariance, we find that (i) the kinetic term of the Abelian 2-form field, and (ii) the FP-ghost term of the Abelian 1-form theory transform in such a manner that they produce a total spacetime derivative******In exactly similar fashion, we note that (i) the kinetic term of the Abelian 1-form gauge field, and (ii) the basic FP-ghost term of the Abelian 2-form theory, transform to a total spacetime derivative under the co-BRST symmetry transformations thereby leading to the invariance of the action integral.. The algebra obeyed by these symmetry operators [cf. Eq. (4)] is reminiscent of the algebra obeyed by the de Rham cohomological operators of differential geometry [cf. Eq. (4)].
One of the highlights of our present investigation is the existence of the discrete duality symmetry transformations in our BRST-quantized theory which has its origin in the Hodge duality operator (on the 3D flat Minkowskian spacetime manifold) as far as the basic gauge fields of the Abelian 1-form and 2-form theories are concerned [cf. Eq. (16) and corresponding footnote]. The interplay between the discrete duality symmetry transformations [cf. Eqs. (3),(3)] and continuous symmetry transformations [cf. Eqs. (3),(3)] for the (co-)BRST invariant Lagrangian density (21) are responsible for the physical realization (29) of the well-known relationship (i.e. $\delta=\pm\,*\,d\,*$) between the (co-)exterior derivatives of differential geometry [14-18]. The most basic relationship (16) between the gauge fields of the Abelian 1-form and 2-form theories establishes clearly that there is a duality symmetry between the free Abelian 2-form and 1-form basic fields††††††We would like to mention that these duality symmetry transformations have not been mentioned in our very recent work [20] on the combined system of the free $3D$ Abelian 1-form and 2-form gauge theories. This is why we have not been able to obtain the analogue of (29). in three (i.e. $D=3$) dimensions of spacetime when they are taken together for specific purpose. It interesting to add that these discrete duality symmetry transformations [cf. Eqs. (3),(3)] are also responsible for establishing a direct connection between the off-shell nilpotent (co-)BRST symmetry transformation operators (cf. Appendix A for details).
We would like to say a few words about the presence of the scalar and pseudo-scalar fields (i.e. $\phi,\,\widetilde{\phi}$) in our (co-)BRST invariant Lagrangian density (21). It is very interesting to observe that (i) the scalar field appears in our theory with a positive kinetic term, and (ii) the pseudo-scalar field is endowed with a negative kinetic term (cf. Sec. 1 for more details. The latter field (having the negative kinetic term) is one of a possible candidates for the “phantom” fields which have become quite popular in the context of the cyclic, bouncing and self-accelerated cosmological models of the Universe (see, e.g. [21-25] and references therein). Such massless fields (having only the negative kinetic term) are also relevant in the context of being one of the passable candidates for dark energy (see, e.g. [26,27] for details) which are supposed to be responsible for the current observations of the accelerated cosmic expansion [32-35] of the Universe.
In our present investigation, we have (i) not computed the Noether conserved charges corresponding to the off-shell nilpotent (co-)BRST symmetry transformations [cf. Eqs. (3),(3)] as well as the bosonic symmetry transformations [cf. Eq. (4)], and (ii) not considered the generalization of the classical gauge-fixed Lagrangian density (19) to its quantum counterpart within the framework of BRST formalism. In his context, it is gratifying to state that, in our very recent paper [36], we have considered the quantum generations of the both the gauge-fixed Lagrangian densities (18) and (19) and obtained the two-to-one mappings between the full set of symmetry transformation operators of the BRST-quantized theory and the set of cohomological operators of differential geometry. In our future endeavors, we plan to devote time to address the above issues and compute the extended BRST algebra among the appropriate set of conserved charges of our theory and establish its deep connection with the algebra obeyed by the de Rham cohomological operators of differential geometry so that we can prove that our present odd dimensional (i.e. $3D$) field-theoretic model of the combined system of the free Abelian 1-form ad 2-form gauge theories is a perfect example for Hodge theory.
Acknowledgment
One of us (RPM) would like to thank E. Harikumar, A. K. Rao , S. K. Panja, Bhagya R. and H. Sreekumar for fruitful discussions on the main topic of our present endeavor.
Appendix A On Direct Connection: Nilpotent $s_{d}$ and $s_{b}$
Besides the algebraic relationship we have obtained between the off-shell nilpotent ($s_{d}^{2}=0,\;s_{b}^{2}=0$) (co-)BRST symmetry transformation operators $(s_{d})s_{b}$ in equation (29), there is a direct relationship between these fermionic (i.e. off-shell nilpotent) symmetry transformation operators due to the presence of the discrete duality symmetry transformation operators [cf. Eqs. (3),(3)]. To corroborate this statement, let us pick-up the nilpotent BRST symmetry transformation: $s_{b}\,B_{\mu\nu}=-\,(\partial_{\mu}C_{\nu}=\partial_{\nu}C_{\mu})$ and apply the discrete duality symmetry transformation operators (3) and (3) on it as follows:
$\big{(}*s_{b}\big{)}\big{(}*B_{\mu\nu}\big{)}=-*\big{(}\partial_{\mu}C_{\nu}-\partial_{\nu}C_{\mu}\big{)}.$ (A.1)
Using the discrete transformations (3) and (3) and taking into account the operation: $*\,s_{b}=s_{d}$, we obtain the following explicit relationship
$s_{d}\Big{(}\pm\,i\,\varepsilon_{\mu\nu\sigma}\,A^{\sigma}\Big{)}=-\,\Big{[}\partial_{\mu}\big{(}\pm\,i\,\bar{C}_{\mu}\big{)}-\partial_{\nu}\big{(}\pm\,i\,\bar{C}_{\mu}\big{)}\Big{]},$ (A.2)
which, finally, leads to the following off-shell nilpotent (i.e. $s_{d}^{2}=0$) co-BRST symmetry transformation for the Abelian 1-form vector field $A_{\mu}$ as:
$s_{d}\,A_{\mu}=-\,\varepsilon_{\mu\nu\sigma}\,\partial^{\nu}\bar{C}^{\sigma}.$ (A.3)
Thus, starting from the BRST symmetry transformation for an antisymmetry tensor gauge field $B_{\mu\nu}$, we have obtained the co-BRST symmetry transformation for the Abelian 1-form vector field $A_{\mu}$ correctly due to the direct application of the discrete duality symmetry transformation operators [cf. Eqs. (3),(3)] and taking into account the discrete duality symmetry transformation for the BRST symmetry transformation operator $s_{b}$ as: $*\,s_{b}=s_{d}$ (which leads to the derivation of the co-BRST symmetry transformation for $A_{\mu}$ field).
Against the backdrop of our above discussions, let us derive the nilpotent BRST symmetry transformation: $s_{b}\,B_{\mu\nu}=-\,(\partial_{\mu}C_{\nu}-\partial_{\nu}C_{\mu})$ from the nilpotent co-BRST symmetry transformation: $s_{d}\,A_{\mu}=-\,\varepsilon_{\mu\nu\sigma}\,\partial^{\nu}\bar{C}^{\sigma}$. Once again, we exploit theoretical strength of the discrete duality symmetry transformations [cf. Eqs. (3),(3)] as follows
$\big{(}*\,s_{d}\big{)}\big{(}*A_{\mu}\big{)}=-\,\varepsilon_{\mu\nu\sigma}\,\partial^{\nu}\big{(}*\,\bar{C}^{\sigma}\big{)},$ (A.4)
where we take: $*\,s_{d}=-\,s_{b}$ as an input. We observe the following from (A.4), namely;
$-\,s_{b}\Big{(}\mp\,\frac{i}{2}\,\varepsilon_{\mu\nu\sigma}\,B^{\nu\sigma}\Big{)}=-\,\varepsilon_{\mu\nu\sigma}\partial^{\nu}\big{(}\pm\,i\,C^{\sigma}\big{)}.$ (A.5)
Simplifying the above relation by exploiting the straightforward property of the $3D$ Levi-Civita tensor, we finally obtain the nilpotent (i.e. $s_{b}^{2}=0$) BRST transformation ($s_{b}$):
$s_{b}\,B_{\mu\nu}=-\,(\partial_{\mu}C_{\nu}-\partial_{\nu}C_{\mu}).$ (A.6)
Thus, we conclude that the off-shell nilpotent ($s_{d}^{2}=0,\;s_{b}^{2}=0$) (co-)BRST symmetry transformation operators $(s_{d})s_{b}$ are directly connected if we use the duality property: $*\,s_{b}=s_{d},\;*\,s_{d}=-\,s_{b}$ and the discrete duality symmetry transformation operators [cf. Eqs. (3),(3)] judiciously. Ultimately, we have been able to establish that the mapping: $s_{b}\Leftrightarrow s_{d}$ can be realized (if we use directly the discrete duality symmetry transformation operators [cf. Eqs. (3),(3)] along with the inputs: $*\,s_{b}=s_{d}$,  $*\,s_{d}=-\,s_{b}$). We have been able to establish these kinds of relationships among all the (co-)BRST symmetry transformation operators $(s_{d})s_{b}$ [cf. Eqs. (3),(3)] as follows:
$s_{b}\,A_{\mu}\;\Leftrightarrow\;s_{d}\,B_{\mu\nu},\qquad\qquad s_{b}\,B_{\mu\nu}\;\Leftrightarrow\;s_{d}\,A_{\mu},\qquad\qquad s_{b}\,C_{\mu}\;\Leftrightarrow\;s_{d}\,\bar{C}_{\mu},$
$s_{b}\,\bar{C}_{\mu}\;\Leftrightarrow\;s_{d}\,C_{\mu},\qquad s_{b}\,\bar{\beta}\;\Leftrightarrow\;s_{d}\,\beta,\qquad s_{b}\,\phi\;\Leftrightarrow\;s_{d}\,\widetilde{\phi},\qquad s_{b}\,\bar{C}\;\Leftrightarrow\;s_{d}\,C,$
$s_{b}\big{[}B_{\mu},\,{\cal B}_{\mu},\,B,\,{\cal B},\,\widetilde{\phi},\,\beta,\,C,\,\rho,\,\lambda\big{]}\;\Leftrightarrow\;s_{d}\big{[}{\cal B}_{\mu},\,B_{\mu},\,{\cal B},\,B,\,\phi,\,\bar{\beta},\,\bar{C},\,\lambda,\,\rho\big{]}.$ (A.7)
It is very interesting to point out that the nilpotent ($s_{b}^{2}=0,\;s_{d}^{2}=0$) BRST ($s_{b}$) and co-BRST ($s_{d}$) symmetry transformation operators are connected with each-other (i.e $s_{b}\;\Leftrightarrow\;s_{d}$) through the fields that are intimately related to each-other by the direct duality symmetry transformations [cf. Eqs. (3),(3)] in our $3D$ field-theoretic model.
We end this Appendix with the following remarks. First, we note that the basic discrete duality symmetry transformations [cf. Eq. (16)] exist between the Abelian 1-form field $A_{\mu}$ and 2-form field $B_{\mu\nu}$ at the level of classically gauge-fixed Lagrangian density [cf. Eq. (12)]. Second, these duality relationships are promoted at the quantum level when the Faddeev-Popov (FP) ghost terms are added to the gauge-fixed Lagrangian density where we have to take the total discrete duality symmetry transformations for the non-ghost fields [cf. Eq. (3)] as well as for the (anti-)ghost fields [cf. Eq. (3)]. Third, we have noted in this Appendix that the total set of discrete duality symmetry transformations [cf. Eqs. (3),(3)] provide a direct mapping between the off-shell nilpotent BRST and co-BRST symmetry transformation operators [cf. Eq. (A)] provided we use: $*\,s_{b}=s_{d}$,  $*\,s_{d}=-\,s_{b}$. Finally, it is very interesting to point out that the duality symmetry transformations between the BRST and co-BRST symmetry operators (i.e. $*\,s_{b}=s_{d}$,  $*\,s_{d}=-\,s_{b}$) are just like the electromagnetic duality symmetry transformations between the electric and magnetic fields of the source-free Maxwell’s theory of electromagnetism in vacuum.